OpenStudy (anonymous):

If f(x) is a continuous function defined for all real numbers, f(-2) = 10, f(-7) = –5, and f(x) = 0 for one and only one value of x, then which of the following could be that x value? a. -3 b. -8 c. 0 d. 2

OpenStudy (anonymous):

@PRAETORIAN.10 @TuringTest @myininaya @AlexandervonHumboldt2 @DanJS @Hero @quickstudent @Here_to_Help15 @HelpBlahBlahBlah @TheSmartOne @satellite73 @Abhisar @Zale101 @AnswerMyQuestions @CausticSyndicalist @emily0824 @iambatman

OpenStudy (zale101):

f(-2) = 10, f(-7) = -5 let's think of them as two given points and we are trying to calculate a continuous function such as a line. f(-2) = 10 the point is (-2,10) f(-7) =-5 the point is (-7,-5) find the equation in slope intercept form and set y equal to 0. Solve for x and there's your answer.

OpenStudy (zale101):

The function is a line because in your question says "f(x) = 0 for one and only one value of x" that pretty much defines a linear function

OpenStudy (solomonzelman):

you have a continuous function \(\large\color{slate}{ f(x) }\), (will assume continuous over the interval \(\large\color{slate}{ (-\infty,\infty)}\) and for real numbers) such that: \(\large\color{slate}{ f(-2) = 10 }\) \(\large\color{slate}{ f(-7) = ~-5 }\) for what value of x can \(\large\color{slate}{ f(x) = 0 }\) ?

OpenStudy (anonymous):

slope is 3

OpenStudy (solomonzelman):

(It is definitely explicit, but) I didn't see a word \(\Large\color{slate}{ \rm linear }\) in there.

OpenStudy (solomonzelman):

I think though, that they want a linear function though.

OpenStudy (solomonzelman):

The slope is 3 that is correct.

OpenStudy (anonymous):

idk wht to do from here

OpenStudy (solomonzelman):

So with a slope 3, and use a point (-2,10) : \(\large\color{slate}{ y-y_1=m(x-x_1)}\)

OpenStudy (solomonzelman):

plug in your information, and re-write in a slope intercept form.

OpenStudy (anonymous):

y=3x+16

OpenStudy (solomonzelman):

\(\large\color{slate}{ y-10=-3(x-(-2))}\) \(\large\color{slate}{ y-10=-3(x+2)}\) \(\large\color{slate}{ y-10=-3x-6}\) .... it is not 16, see the err ?

OpenStudy (anonymous):

there isnt a negative in front of the 3 for the slope right?

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

oh, true that

OpenStudy (solomonzelman):

SO you are correct then, it is \(\large\color{slate}{ y=3x+16}\)

OpenStudy (anonymous):

so now wht

OpenStudy (solomonzelman):

\(\large\color{slate}{ f(x)=0}\) would be an x intercept. Set y=0 equal zero, and solve for x. \(\large\color{slate}{ 0=3x+16}\)

OpenStudy (anonymous):

-5.333333=x

OpenStudy (anonymous):

and repeating

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

yes, -16/3

OpenStudy (solomonzelman):

I would cite the exact value.

OpenStudy (anonymous):

now what

OpenStudy (solomonzelman):

But this again, is assuming that your function is linear.

OpenStudy (anonymous):

If f(x) is a continuous function defined for all real numbers, f(-2) = 10, f(-7) = -5, and f(x) = 0 for one and only one value of x, then which of the following could be that x value? a. -3 b. -8 c. 0 d. 2

OpenStudy (solomonzelman):

not an option, so it is not (necessarily) a linear function.

OpenStudy (anonymous):

How do you mean, it's not one of the answer choice provided?

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

I mean that -16/3 isn't an option.

OpenStudy (anonymous):

yes ik tht... so there are more steps, correct?

OpenStudy (anonymous):

ohh okay

OpenStudy (solomonzelman):

if you have points, (-2,10) (-7,-5) and only think about f(x) is given that it is continuous, but not necessarily is linear, then,

OpenStudy (solomonzelman):

it can be (technically any point. Continuous means that it is explicit (not more than one x value for a y value) and that it can be drawing without picking the pencil of the paper.